OpenStudy (anonymous):

Suppose p varies directly as d, and p = 2 when d = 7. What is the value of d when p = 10? A. d= 7/5 B. d= 15 C. d= 20/7 D. d= 35

OpenStudy (stamp):

p / d = 2 / 7 7 p = 2 d d = 7p / 2 d(p) = 7p / 2 Find d when p is 10. d(10) = 7(10) / 2

OpenStudy (stamp):

d(10) = 70 / 2

OpenStudy (anonymous):

70/2 isnt one of the choices

OpenStudy (stamp):

@Cookie_009 you realize that math is a cumulative subject, correct?

OpenStudy (stamp):

You cannot just learn one thing, forget it, and then hope to succeed next year

OpenStudy (stamp):

Do you remember how to divide by 2?

OpenStudy (anonymous):

yaa

OpenStudy (anonymous):

hellllooo

OpenStudy (stamp):

70 / 2 =

OpenStudy (anonymous):

i got it nevermind its d